OpenStudy (anonymous):

find the equation of the line having the given slope and containing the given point m=7/8(7,-5)

OpenStudy (anonymous):

please help

OpenStudy (anonymous):

Use the point-gradient formula for a straight line: \[y-y_1=m(x-x_1)\]So here,\[y-(-5)=\frac{7}{8}(x-7)\]\[y+5=\frac{7}{8}(x-7)\]Simplify.

OpenStudy (anonymous):

Multiply both sides by 8, and expand the right-hand side.\[8y+40=7x-49\]

OpenStudy (anonymous):

is that the final answer

OpenStudy (anonymous):

\[7x-8y-89=0\]

OpenStudy (anonymous):

That's the equation in general form.

OpenStudy (anonymous):

darn.....i need a straight forward answer

OpenStudy (anonymous):

Well, that's pretty much it. I don't know what you're looking for. You asked for the equation of the line with given slope containing a given point...that's the answer. If you need it in point-gradient form, you just need to rearrange it in the form y=...

OpenStudy (anonymous):

well thank yoou very very much medal for u

OpenStudy (anonymous):

np :)